A mathematics placement test is given to all entering freshmen at Eastern

University. n = 55,  ̄x = 4, ̄y = 12, "\sum"ni=1 xi2=232, "\sum"ni=1 xiyi=318 .Fit a simple

linear regression model between x and y


1
Expert's answer
2021-09-27T03:15:09-0400

Regression equation: "y=a+bx."

"b=\frac{\Sigma xy}{\Sigma x^2}=
\frac{318}{232}=1.3707."

"a=\bar y-b\bar x=12-1.3707*4=6.5172."

Thus, the regression equation is "y=6.5172+1.3797x."